# On Entropy Control in LLM-RL Algorithms

### v2

@@ 1 Introduction
The methods used in LLM-RL are predominantly policy-gradient based, e.g., the PPO (Schulman et al., 2017) family. Policy gradient based methods reinforce the sampled actions that lead to higher rewards compared to other sampled actions. However, when the optimal actions are not sampled, the policy gradient methods can over-reinforce the sampled locally optimal actions, ultimately resulting in the policy stuck at suboptimal points (Agarwal et al., 2021). The sub-optimal actions can be meaningless in deep RL and oftentimes have a large performance gap from the optimal ones (Henderson et al., 2018), e.g., in LLM tasks, the policy can be stuck at producing the correct format but incorrect results. A straightforward remedy for the issue was the so-called entropy-regularized RL methods (Williams & Peng, 1991), where the policy maximizes a sum of rewards and some entropy bonus (policy randomness). This technique was commonly used in policy-gradient methods including A3C (Mnih et al., 2016), PPO (Schulman et al., 2017) and SAC (Haarnoja et al., 2018), providing strong benefits in tasks requiring hierarchical behaviors. Intuitively, the entropy bonus keeps the policy random and explorative, thus prevents the policy from over-reinforcing certain actions and getting stuck. Moreover, entropy regularization is shown to provide strong optimization benefits both empirically (Ahmed et al., 2019) and theoretically (Mei et al., 2020; Klein et al., 2023).
@@ 1.1 Related works
Entropy regularization in RL. Entropy-regularized RL was initially introduced in (Williams & Peng, 1991). It has been commonly used in various popular policy-based deep RL algorithms (Mnih et al., 2016; Schulman et al., 2017; Haarnoja et al., 2018) which have provided ample empirical evidence for its effectiveness in robotic and games tasks. Entropy regularization’s optimization benefits have also been empirically (Ahmed et al., 2019) and theoretically (Mei et al., 2020; Klein et al., 2023) studied. However, it does not give notable performance gains for LLMs (see, e.g., Section 5 and (Cui et al., 2025)). As a result, alternative entropy control techniques are often adopted. In (Zhang et al., 2024) reshapes the reward function to regulate the policy. Or in a concurrent work (Cui et al., 2025), the algorithm clips or regulates the parts the policy update that decrease entropy too much. To our best knowledge, existing works do not answer the question of why and when entropy regularization can fail in LLM-RL, and have not uncovered the potential benefits of entropy bonus.
@@ 2 Preliminaries
In this section, we will first give formal definitions of some RL concepts, and then introduce several prominent existing policy optimization algorithms.
@@
where $\hat{A}_{t}$ is an estimate of the advantage function. GRPO uses a trajectory-level Monte-Carlo estimate of the advantage. DAPO additionally decouples the clip ratio by setting different $\epsilon_{\rm low},\epsilon_{\rm high}$ and incorporates extra sampling constraints and overlong response penalty. Given a suitable clip range, the PPO-clip algorithm can be viewed as a policy gradient algorithm (Jin et al., 2023).
@@ 3 A theory on entropy effect in policy gradient based LLM-RL
where $\theta_{s,a}$ is the logit of token $a$ given input $s$. The policy gradient based algorithms without entropy regularization, including PPO-clip, are generally guaranteed to converge to an $\epsilon$-stationary point of the RL objective $V^{\pi_{\theta}}({\mathcal{D}})$ satisfying $\|\nabla V^{\pi_{\theta}}({\mathcal{D}})\|\leq\epsilon$ (Agarwal et al., 2021; Jin et al., 2023). When doing policy optimization without regularization, Cui et al. (2025) observes that the policy entropy quickly diminishes as performance increases, and ultimately performance saturates when entropy completely collapses. In the following result, we give some theoretical insights into this case.
@@
    $V^{\pi^{*}}(s_{0})-V^{\pi_{\theta}}(s_{0})\leq\frac{1}{C^{\pi_{\theta}}(s_{0})}\epsilon$
@@
The first bullet (I) suggests the policy entropy is an indicator of the policy stationarity, that is, a small entropy indicates a small policy gradient $\|\nabla V^{\pi_{\theta}}({\mathcal{D}})\big\|$ and the convergence of the policy. The second bullet (II) quantifies the actual performance of the almost stationary policy, where the reward on query $s_{0}$ is bounded by $\mathcal{O}(\epsilon/C^{\pi_{\theta}}(s_{0}))$. The factor $C^{\pi_{\theta}}(s_{0})$ can be controlled (bounded away from 0) when the initial LLM and the RL algorithm can sufficiently explore the optimal response to $s_{0}$. For example, one can either use a large batch size (Klein et al., 2023) or a strong initial model (Weissmann et al., 2024) to control $C^{\pi_{\theta}}(s_{0})$. In this case, the performance is ultimately bounded by $\mathcal{O}(\epsilon)$. The error $\epsilon$ decreases with prolonged RL training, while it usually cannot decrease to 0 due to the presence of sampling noise or the advantage estimation error.
@@
On the other hand, the maximum entropy RL optimizes the entropy-regularized reward sum $V_{\lambda}^{\pi_{\theta}}({\mathcal{D}})$ (Williams & Peng, 1991). In non-LLM deep RL tasks, this method has long been popular and can significantly outperform methods without entropy control (Mnih et al., 2016; Haarnoja et al., 2018). However, experiments (to be shown in Section 5) show that traditional entropy regularization gives weak to no gains in LLM-RL training. In the next result, we give theoretical insight into this issue.
@@
$V^{\pi^{*}}(s_{0})-V^{\pi_{\theta}}(s_{0})\leq\frac{1}{C_{\lambda}^{\pi_{\theta}}(s_{0})}\frac{\epsilon^{2}}{2\lambda}+\lambda H\log\frac{|{\mathcal{A}}|}{|{\mathcal{A}}_{H}^{*}(s_{0})|^{\frac{1}{H}}}$
@@
Similar conditions to Propositions 1.(II)&2 have been derived in (Mei et al., 2020) for the discounted infinite horizon MDPs, while our results hold for the finite horizon MDPs under a concatenation transition. Proposition 2 also provides a more accurate bound for the entropy bias.
Entropy regularization suffers from immense response space with sparse optimality in LLM tasks. As compared to no entropy control case in Proposition (II), the above bound’s dependence on $\epsilon$ improves to $\mathcal{O}(\epsilon^{2}/2\lambda)$ when $C_{\lambda}^{\pi_{\theta}}(s_{0})$ is bounded away from 0. However, this optimization benefit does not come free as a bias term is introduced. The entropy bias is $\mathcal{O}(H\log(|{\mathcal{A}}|/|{\mathcal{A}}_{H}^{*}(s_{0})|^{\frac{1}{H}}))$, which increases with the response space size $H\log|{\mathcal{A}}|$ and the sparsity of optimal responses $\log(1/|{\mathcal{A}}_{H}^{*}(s_{0})|)$. The bias is especially ubiquitous in LLM-RL, where the response space is typically extremely large (hundreds of thousands tokens to sample from in each step) as compared to that in, e.g., classic control and games where the action space size and horizon are typically at the hundreds (Brockman et al., 2016; Silver et al., 2017). To better demonstrate this effect, we report a numerical test result in Figure 1. The test is done in a synthetic MDP with $|{\mathcal{A}}|=10^{5}$, and the optimal action is sparse with various numbers in $\{1,5,10,15\}$ (see Appendix A.5 for details). It is observed that entropy regularization leads to gains over no-regularization when number of optimal actions is $10,15$, but offers no gains when the optimal action becomes too sparse with fewer than $5$ optimal actions.
@@ 5.1 Training details
Models, training datasets and baselines. The algorithms are tested in two training settings: (a) we train the Qwen2.5-math-1.5b base model on the MATH dataset (Hendrycks et al., 2021), which contains $7500$ math problems with various difficulties and covers multiple mathematical areas; (b) we also train the DeepSeek-R1-distilled-Qwen-1.5b (DeepSeek-AI, 2025) model on 4040k verifiable queries from the OpenR1-math (Open-R1, 2025) dataset, which is derived from Numina-math dataset (Li et al., 2024). We compare our algorithm with GRPO and the conventional entropy regularization method which we call EntReg, where the GRPO objective is augmented with the original entropy bonus used in (Mnih et al., 2016; Schulman et al., 2017).
@@
Hyper-parameter settings. The tests are based on the verl framework (Sheng et al., 2025) [^1] The code will be released as soon as an internal mandatory pre-open-source review is completed.. For experiment setting (a), we use AdamW optimizer with a learning rate of $2\times 10^{-6}$. We set the max response length as $3072$. We use a batch size of $512$, and for each query we roll out $16$ responses. For AEnt, we use the GRPO objective as ${\mathcal{L}}_{\rm PO}$. We use a clamping percentage $p=0.33$, and set $\tilde{{\mathcal{H}}}_{\rm low}\!=\!0.15$ and $\tilde{{\mathcal{H}}}_{\rm high}\!=\!0.24$. We use an initial entropy coefficient $\lambda=0.002$, and start updating it from the third epoch with $\beta=0.002$. We clip the coefficient in between $0.0006$ and $0.009$. For EntReg method, we use the traditional entropy bonus with a fixed entropy coefficient of $0.002$. For experiment setting (b), we use a learning rate of $1\!\times\!10^{-6}$, a max response length of $7168$, a batch size of $256$ and for each query we roll out $8$ responses. We use $p=0.25$, $\tilde{{\mathcal{H}}}_{\rm low}\!=\!0.35$ and $\tilde{{\mathcal{H}}}_{\rm high}\!=\!0.62$, an initial $\lambda=3\!\times\!10^{-4}$, and start updating it from the second epoch with $\beta=10^{-4}$. We clip the coefficient in between $4\times 10^{-5}$ and $0.001$.
@@ 5.2 Performance analysis
An observation on the test score and the entropy trend. An interesting observation from Figures 4(a) is that after around $175$ steps (collapse time), the policy entropy of GRPO largely depletes and the entropy of EntReg starts to drastically fluctuate, while AEnt’s policy entropy is kept stable. Then one can observe from Figure 3(a) that the test score of GRPO and EntReg plateaus around the same step, while the score of AEnt continues to improve and surpasses the baselines past the collapse time. This observation is consistent with our intuition and theoretical analysis: after GRPO’s entropy collapse, its policy becomes concentrated on few paths and no new information can be gained in the sampling process, ultimately leading to the stagnancy of the learning process. This is predicted by Proposition (I) that the policy will become stationary after entropy depletion.
@@ 5.4 Time complexity
6 Conclusion
In this work, we showed that entropy regularization suffers from large bias in LLM-RL training. As a remedy of this issue, we propose an entropy control method that utilizes a clamped entropy bonus with an automatically adjusted coefficient. We show that AEnt consistently outperforms competitive baselines across multiple benchmarks. We believe AEnt can demonstrate more significant advantages if tested on larger models with more compute. In this work, we did not include a theoretical analysis of the clamped entropy. In addition, we believe the choice of the clamped space ${\mathcal{A}}(s)$ is crucial to the algorithm’s effectiveness, and finding a better choice can potentially yield significant performance gains. We leave these studies for future work.

### v3

@@ 1 Introduction
The methods used in LLM-RL are predominantly policy-gradient based, e.g., the PPO (Schulman et al., 2017) family. Policy gradient based methods reinforce the sampled actions that lead to higher rewards compared to other sampled actions. However, when the optimal actions are not sampled, the policy gradient methods can over-reinforce the sampled locally optimal actions, ultimately resulting in the policy stuck at suboptimal points (Agarwal et al., 2021). The sub-optimal actions can be meaningless in deep RL and oftentimes have a large performance gap from the optimal ones (Henderson et al., 2018), e.g., in LLM tasks, the policy can be stuck at producing the correct format but incorrect results. A straightforward remedy for the issue was the so-called entropy-regularized RL methods (Williams and Peng, 1991), where the policy maximizes a sum of rewards and some entropy bonus (policy randomness). This technique was commonly used in policy-gradient methods including A3C (Mnih et al., 2016), PPO (Schulman et al., 2017) and SAC (Haarnoja et al., 2018), providing strong benefits in tasks requiring hierarchical behaviors. Intuitively, the entropy bonus keeps the policy random and explorative, thus prevents the policy from over-reinforcing certain actions and getting stuck. Moreover, entropy regularization is shown to provide strong optimization benefits both empirically (Ahmed et al., 2019) and theoretically (Mei et al., 2020; Klein et al., 2023).
@@ 1.1 Related works
Entropy regularization in RL. Entropy-regularized RL was initially introduced in (Williams and Peng, 1991). It has been commonly used in various popular policy-based deep RL algorithms (Mnih et al., 2016; Schulman et al., 2017; Haarnoja et al., 2018) which have provided ample empirical evidence for its effectiveness in robotic and games tasks. Entropy regularization’s optimization benefits have also been empirically (Ahmed et al., 2019) and theoretically (Mei et al., 2020; Klein et al., 2023) studied. However, it does not give notable performance gains for LLMs (see, e.g., Section 5 and (Cui et al., 2025)). As a result, alternative entropy control techniques are often adopted. In (Zhang et al., 2024) reshapes the reward function to regulate the policy. Or in a concurrent work (Cui et al., 2025), the algorithm clips or regulates the parts the policy update that decrease entropy too much. To our best knowledge, existing works do not answer the question of why and when entropy regularization can fail in LLM-RL, and have not uncovered the potential benefits of entropy bonus.
@@ 2 Preliminaries
In this section, we will first give formal definitions of some RL concepts, and then introduce several prominent policy optimization algorithms.
@@
where $\hat{A}_{t}$ is an estimate of the advantage function. GRPO uses a Monte-Carlo estimate of the trajectory-level advantage. DAPO additionally decouples the clip ratio by setting different $\epsilon_{\rm low},\epsilon_{\rm high}$ and incorporates extra sampling constraints and overlong response penalty. Given a suitable clip range, the PPO-clip algorithm can be viewed as a policy gradient algorithm (Jin et al., 2023).
@@ 3 A theory on entropy effect in policy gradient based LLM-RL
where $\theta_{s,a}$ is the logit of token $a$ given input $s$. The policy gradient based algorithms without entropy regularization, including PPO-clip, are generally guaranteed to converge to an $\epsilon$-stationary point of the RL objective $V^{\pi_{\theta}}({\mathcal{D}})$ satisfying $\|\nabla V^{\pi_{\theta}}({\mathcal{D}})\|\leq\epsilon$ (Agarwal et al., 2021; Jin et al., 2023). When doing policy optimization without regularization, (Cui et al., 2025) observes that the policy entropy quickly diminishes as performance increases, and ultimately performance saturates when entropy completely collapses. In the following result, we give some theoretical insights into this observation.
@@
    $V^{\pi^{*}}(s_{0})-V^{\pi_{\theta}}(s_{0})\leq\frac{1}{C^{\pi_{\theta}}(s_{0})}\epsilon$ (3.1)
@@
The first bullet (I) suggests the policy entropy is an indicator of the policy stationarity, that is, a small entropy indicates a small policy gradient $\|\nabla V^{\pi_{\theta}}({\mathcal{D}})\big\|$ and the convergence of the policy. The second bullet 3.1 quantifies the actual performance of the almost stationary policy, where the reward optimality gap on query $s_{0}$ is bounded by $\mathcal{O}(\epsilon/C^{\pi_{\theta}}(s_{0}))$. The factor $C^{\pi_{\theta}}(s_{0})$ can be controlled (bounded away from 0) when the initial LLM and the RL algorithm can sufficiently explore the optimal response to $s_{0}$. For example, one can either use a large batch size (Klein et al., 2023) or a strong initial model (Weissmann et al., 2024) to control $C^{\pi_{\theta}}(s_{0})$. In this case, the performance is ultimately bounded by $\mathcal{O}(\epsilon)$. The error $\epsilon$ decreases with prolonged RL training, while it usually cannot decrease to 0 due to the presence of sampling noise or the advantage estimation error.
@@
On the other hand, the maximum entropy RL optimizes the entropy-regularized reward sum $V_{\lambda}^{\pi_{\theta}}({\mathcal{D}})$ (Williams and Peng, 1991). In non-LLM deep RL tasks, this method has long been popular and can significantly outperform methods without entropy control (Mnih et al., 2016; Haarnoja et al., 2018). However, experiments (to be shown in Section 5) show that traditional entropy regularization gives weak to no gains in LLM-RL training. In the next result, we give theoretical insight into this issue.
@@
$V^{\pi^{*}}(s_{0})-V^{\pi_{\theta}}(s_{0})\leq\frac{1}{C_{\lambda}^{\pi_{\theta}}(s_{0})}\frac{\epsilon^{2}}{2\lambda}+\lambda H\log\frac{|{\mathcal{A}}|}{|{\mathcal{A}}_{H}^{*}(s_{0})|^{\frac{1}{H}}}$ (3.2)
@@
Similar conditions to Propositions 1.3.1&2 have been derived in (Mei et al., 2020) for the discounted infinite horizon MDPs, while our results hold for the finite horizon MDPs under a concatenation transition. Proposition 2 also provides a more accurate bound for the entropy bias.
Entropy regularization suffers from immense response space with sparse optimality in LLM tasks. As compared to no entropy control case in Proposition 3.1, the above bound’s dependence on $\epsilon$ improves to $\mathcal{O}(\epsilon^{2}/2\lambda)$ when $C_{\lambda}^{\pi_{\theta}}(s_{0})$ is bounded away from 0. However, this optimization benefit does not come free as a bias term is introduced. The entropy bias is $\mathcal{O}(H\log(|{\mathcal{A}}|/|{\mathcal{A}}_{H}^{*}(s_{0})|^{\frac{1}{H}}))$, which increases with the response space size $H\log|{\mathcal{A}}|$ and the sparsity of optimal responses $\log(1/|{\mathcal{A}}_{H}^{*}(s_{0})|)$. The bias is especially ubiquitous in LLM-RL, where the response space is typically extremely large (hundreds of thousands tokens to sample from in each step) as compared to that in, e.g., classic control and games where the action space size and horizon are typically at the hundreds (Brockman et al., 2016; Silver et al., 2017). To better demonstrate this effect, we report a numerical test result in Figure 1. The test is done in a synthetic MDP with $|{\mathcal{A}}|=10^{5}$, and the optimal action is sparse with various numbers in $\{1,5,10,15\}$ (see Appendix A.5 for details). It is observed that entropy regularization leads to gains over no-regularization when number of optimal actions is $10,15$, but offers no gains when the optimal action becomes too sparse with fewer than $5$ optimal actions.
@@ 5.1 Training details
Models, training datasets and baselines. The algorithms are tested in multiple training settings: (a) we train the Qwen2.5-math-1.5b base model on the MATH dataset (Hendrycks et al., 2021), which contains $7500$ math problems with various difficulties and covers multiple mathematical areas; (b) we train the DeepSeek-R1-distilled-Qwen-1.5b (DeepSeek-AI, 2025) model on 4040k verifiable queries from the OpenR1-math (Open-R1, 2025) dataset, which is derived from Numina-math dataset (Li et al., 2024). In addition, we also train the Qwen2.5-math-7b model on 6k samples from DeepMath dataset (He et al., 2025), the results of which is deferred to Appendix A.6. We compare our algorithm with GRPO and the conventional entropy regularization method which we call EntReg, where the GRPO objective is augmented with the original entropy bonus used in (Mnih et al., 2016; Schulman et al., 2017).
@@
Hyper-parameter settings. The tests are based on the verl framework (Sheng et al., 2025).  [^1] Our code is available at https://github.com/antgroup/AEnt. When training Qwen2.5-math-1.5b base model on the MATH dataset, we use AdamW optimizer with a learning rate of $2\times 10^{-6}$. We set the max response length as $3072$. We use a batch size of $512$, and for each query we roll out $16$ responses with default sampling parameters (top-p and temperature set as $1$). For AEnt, we use the GRPO objective as ${\mathcal{L}}_{\rm PO}$. We use a clamping percentage $p=0.33$, and set $\tilde{{\mathcal{H}}}_{\rm low}\!=\!0.15$ and $\tilde{{\mathcal{H}}}_{\rm high}\!=\!0.24$. We use an initial entropy coefficient of $0.002$, and start updating the coefficient from the third epoch with $\beta=0.002$. We clip the coefficient in between $0.0006$ and $0.009$. For EntReg method, we use the traditional entropy bonus with a fixed entropy coefficient of $0.002$. When training DeepSeek-R1-distilled-Qwen model on the OpenR1-math dataset, we use a learning rate of $1\!\times\!10^{-6}$, a max response length of $7168$, a batch size of $256$ and for each query we roll out $8$ responses. We use $p=0.25$, $\tilde{{\mathcal{H}}}_{\rm low}\!=\!0.35$ and $\tilde{{\mathcal{H}}}_{\rm high}\!=\!0.62$, an initial coefficient of $3\!\times\!10^{-4}$, and start updating the coefficient from the second epoch with $\beta=10^{-4}$. We clip the coefficient in between $4\times 10^{-5}$ and $0.001$.
@@ 5.2 Performance analysis
An observation on the test score and the entropy trend. An interesting observation from Figures 4(a) is that after around $175$ steps (collapse time), the policy entropy of GRPO largely depletes and the entropy of EntReg starts to drastically fluctuate, while AEnt’s policy entropy is kept stable. Then one can observe from Figure 3(a) that the test score of GRPO and EntReg plateaus around the same step, while the score of AEnt continues to improve and surpasses the baselines past the collapse time. This observation is consistent with our intuition and theoretical analysis: after GRPO’s entropy collapse, its policy becomes concentrated on few paths and no new information can be gained in the sampling process, ultimately leading to the stagnancy of the learning process. This is predicted by Proposition (I) that the policy will become stationary after entropy depletion. Additionally, it can be observed from Figure 4 that the entropy regularization methods result in slightly longer response in the mid/end of the training period. The potential reason is that the regularization makes the model less certain, and thus the models tend to continue its generation, resulting in longer response. Nonetheless, the increase in response length is relatively mild and we did not observe a major drawback in reasoning efficiency.
@@ 5.4 Time complexity
6 Conclusion and future directions
In this work, we showed that entropy regularization suffers from large bias in LLM-RL training. As a remedy of this issue, we propose an entropy control method that utilizes a clamped entropy bonus with an automatically adjusted coefficient. We show that AEnt consistently outperforms competitive baselines across multiple benchmarks. We believe AEnt can demonstrate more significant advantages if tested on larger models with more compute. In this work, we did not include a theoretical analysis of the clamped entropy. In addition, we believe the choice of the clamped space ${\mathcal{A}}(s)$ is crucial to the algorithm’s effectiveness, and finding a better choice can potentially yield significant performance gains. For example, one may consider removing actions that are redundant (Baram et al., 2021; Zhong et al., 2024) or grouping similar actions in entropy calculation. Alternatively, since computing entropy on token level suffers from the large dimensionality of LLM vocabulary, it can be beneficial to design entropy regularizer in the state or action representation space (Tennenholtz and Mannor, 2019; Tavakoli et al., 2018). We leave these studies to future works.
@@ A.5 Toy verification in Figure 1
A.6 Additional experiments
In this section, we report some additional experimental results.
Experimental details. The algorithms are tested on Qwen2.5-math-7b base model on $6144$ samples from the DeepMath dataset (He et al., 2025). We randomly select queries with over-long filtering (no more than $1024$ tokens) from the dataset. The evaluation method is consistent with that in Section 5.1. We use AdamW optimizer with a learning rate of $1\times 10^{-5}$. We set the max response length as $3072$. We use a batch size of $512$, and for each query we roll out $8$ responses. For AEnt, we use the GRPO objective as ${\mathcal{L}}_{\rm PO}$. We use a clamping percentage $p=0.33$, and set $\tilde{{\mathcal{H}}}_{\rm low}\!=\!0$ and $\tilde{{\mathcal{H}}}_{\rm high}\!=\!0.3$. We use an initial entropy coefficient of $0.002$, and start updating the coefficient from the third epoch with $\beta=0.001$. We clip the coefficient in between $0.0006$ and $0.005$. For EntReg method, we use the traditional entropy bonus with an entropy coefficient of $0.002$.
Figure 7: Results of training Qwen2.5-Math-7B on 6k samples from the DeepMath dataset.
MATH-Hard MATH-500 AIME24 Minerva Olympiad AMC
Base 0.443 0.626 0.183 0.143 0.290 0.469
GRPO 0.622 0.808 0.250 0.358 0.412 0.631
EntReg 0.620 0.810 0.214 0.365 0.437 0.644
AEnt 0.657 0.828 0.258 0.379 0.493 0.637
Table 3: Benchmark scores of training Qwen2.5-Math-7B on 6k samples from the DeepMath dataset. Bold numbers indicate the best result on the benchmark.
Observations. The results are reported in Figure 7 and Table 3. The performance observation of Figure is overall consistent with that in Section 5.2. It can be observed that AEnt is able to outperform the baselines (left plot) potentially by preventing an early entropy collapsing (middle plot). The response length of the algorithms is overall similar towards the late training period. AEnt’s training reasoning efficiency is thus similar to the baselines in this test.